OpenStudy (anonymous):

What is the domain of the function? y=(√5x-10) What is the range of the function? y=(√x+2)-6 Please explain!!

OpenStudy (anonymous):

is it \[f(x)=\sqrt{5x-10}\]?

OpenStudy (anonymous):

The question said y=

OpenStudy (anonymous):

yeah, because it was written by an idiot, this beginning the life long confusion between a function and an equation no problem we can still solve if

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

you cannot take the square root of a negative number, so your job here is to solve \[5x-10\geq 0\] for \(x\) which takes only two steps

OpenStudy (anonymous):

add 10, divide by 5 and get \[5x-10\geq 0\\ 5x\geq 10\\ x\geq 2\] is your domain

OpenStudy (anonymous):

Ok that makes sense

OpenStudy (anonymous):

for the second one, is it \[y=\sqrt{x+2}-6\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

in this case you are looking for the possible \(y\) values, not the \(x\) values it is always the case that \(\sqrt{\text{whatever}}\geq 0\) and so \[\sqrt{\text{whatever}}-6\geq -6\] your range is \(y\geq -6\)

OpenStudy (anonymous):

Ok, Thanks! Your explanations made alot of sense!

OpenStudy (anonymous):

yw, glad to help

OpenStudy (anonymous):

now don't tell your math teacher, but \(y=expression\) is not a function, it is an equation also don't tell him or her that you do not "find" the domain of a function, you should be told the domain of a function; it is part of the definition we will keep that between us

OpenStudy (anonymous):

Ok :)